S=-500+0.36y

I=8000

a) consumption function

b) equilibrium level of national income

c) multiplier





1
Expert's answer
2021-05-24T13:20:15-0400

a) The consumption function

"C=500+0.64y"



b)The equilibrium level of national income.

"Y=500+0.64y +8000"

"Y=8500+0.64y"

"0.36y=8500"

"Y=\frac{8500}{0.36}"

"Y=23611.11"



c)The Multiplier

"\frac{1}{1-mpc}"

"\frac{1}{1-0.64}"

"K=2.78"